i want to know how you find the area of an equilateral triangle (regular 3-gon) with the given measurement.
6-inch side.
A = sq. in.
: i want to know how you find the area of an equilateral triangle (regular 3-gon) with the given measurement.
6-inch side.
A = sq. in.

Answer by Earlsdon(3748)

You can use Heron's formula for finding the area of any triangle when you know only the lengths of the three sides.
Heron's formula is:
A = sqrt(s(s-a)(s-b)(s-c)) where s = the semi-perimeter of the triangle and a, b, and c are the lengths of the three sides.
In your problem, s = (6+6+6)/2 = 18/2 = 9 and a, b, and c = 6 inches each.
A = sqrt(9(9-6)(9-6)(9-6))
A = sqrt(9*3*3*3)
A = sqrt(9*27)
A = sqrt(243)
A = 15.6sq.in. to the nearest tenth.